Assume you are managing a food processing plant in Ethiopia. The demand function

for one of your product is given as Qd=50-2p.

a) Find the point price elasticity if price is 15 ETB? Is it elastic or inelastic?

b) How do you interpret the elasticity result?

c) In order to get more revenue what will be your recommendation. Is it to increase

price or decrease price? Why?

d) Describe at least four determinants of the price elasticity demand for the food

product?


1
Expert's answer
2021-06-13T17:38:57-0400

(a) Let's first find the quantity of the product if price is 15 ETB:


"Q_d=50-2\cdot15=20."


Let's also write the inverse demand function:


"P=25-0.5Q_d."


We can find the point elasticity of demand as follows:


"E_d=\dfrac{-1}{slope}\cdot\dfrac{P}{Q_d},""E_d=\dfrac{-1}{-0.5}\cdot\dfrac{15}{20}=1.5"

(b) As we can see from calculations, at this point the price elasticity of demand is elastic

("E_d>1"). Therefore, there is the large change in quantity demanded due to the change in price.

(c) In order to get more revenue we need to decrease the price, because the pecrentage decrease in price will result in larger percentage increase in the quantity of product being sold, therefore, the total revenue will increase.

(d) There are the following determinants of the price elasticity demand for the food

product: substitute products, proportion of income, luxury or necessity, time to respond.